OpenStudy (quickstudent):

In a Hardy-Weinberg problem, the 2pq represents: A. homozygous dominants B. heterozygotes C. homozygous recessives D. hemizygous individuals

OpenStudy (quickstudent):

@N.atia?

OpenStudy (anonymous):

im sorry i still didnt take the hardy weinberg principle

OpenStudy (quickstudent):

@Whitemonsterbunny17 ?

OpenStudy (abhisar):

Hello @quickstudent !

OpenStudy (abhisar):

In Hardy Weinberg Law i.e \(\large\sf P^2+Q^2+2PQ\), P\(^2\) = Frequency of homozygous dominant Q\(^2\) = Frequency of homozygous recessive 2PQ = Frequency of heterozygous.

OpenStudy (abhisar):

Got it ?

OpenStudy (nincompoop):

hardy-weinberg is just a simple algebra of a quadratic equation \[(a+b)^2 = a^2+2ab+b^2 \] where P is homo dominant and Q is homo recessive, and of course PQ would be homo and homo so that's a hetero whatever this is, it has no real-world application, though the point of knowing this is just about getting an understanding of distribution of traits under controlled circumstances. NOW, go forth and read your biology/genetics book and learn some more instead of relying on other people :)